# Sufficient conditions for QMC analysis of finite elements for

### v1

@@ 3.1 The setting
If $J$ is a continuous linear functional then Assumption (A4) holds with $r=1$, since we have $|\tfrac{\partial}{\partial\boldsymbol{q}_{h}}J(\boldsymbol{q}_{h})|\leq C_{J}$ for all $\boldsymbol{q}_{h}$ some constant $C_{J}$, and the chain rule dictates that
@@
which is an even stronger estimate than (A4) with $r=1$.
@@
then Assumption (A4) holds. This can be seen using the Leibniz rule for the product of two functions $f({\boldsymbol{{y}}})$ and $g({\boldsymbol{{y}}})$, whereby
@@
where we used the Cauchy–Schwarz inequality twice. This verifies (A4) with $r=2$.
The next important assumption is on the random parameter. This assumption is motivated by [3, 4]. By Pringsheim’s theorem (cf., e.g., [2, p. 169]), the case $\sigma=1$ in (16) below corresponds to the class of holomorphic functions; the cases of $\sigma\in[0,1)$ may therefore be viewed in some abstract way as slightly smoother than holomorphic functions. Since $\sigma=1$ already recovers the optimal rate, estimating $1\leq(|\boldsymbol{\nu}|!)^{\sigma}\leq|\boldsymbol{\nu}|!$ for $0\leq\sigma<1$ is effectively lossless. Thus, we avoid including the case $\sigma\in[0,1)$ in the analysis since some later inequalities exploit the fact that $x\mapsto x^{\sigma}$, $\sigma\geq 1$, is convex for $x\geq 0$. The assumed cases of $\sigma\geq 1$ covers holomorphic and less regular parametric dependence.
  1. (A5)
@@
$a(\boldsymbol{x},\boldsymbol{y})=\exp\left(\sum_{j\in{\mathbb{N}}}\frac{\xi(y_{j})}{j^{2}}\right).$
If we choose $\xi$ as the identity, we obtain the (classical, analytic) lognormal case, and for the (non-analytic) lognormal Gevrey-$\sigma$ case, we choose
@@
To complete this subsection, we point out a consequence of Assumption (A2), (A3) and (A5) (namely Lemma 3.5 below) that we shall use in our analysis. We also recall two other results we need whose proofs can be found elsewhere.
Lemma 3.4 (see [21, Lemma 3.1]).
@@
Conditions (A2), (A3) and (A5) imply
(17) $\|\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{Q_{h}}\leq C_{S}\|a({\boldsymbol{{y}}})\|^{1/2}_{L^{\infty}(D)}\|f\|_{L^{2}(D)}(C_{R}C_{G}+1)^{|\boldsymbol{\nu}|}(|\boldsymbol{\nu}|!)^{\sigma}\boldsymbol{b}^{\boldsymbol{\nu}}.$
Proof 3.6.
Set $\Upsilon_{\boldsymbol{\nu}}=\|\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{Q_{h}}$. Then
$\Upsilon_{{\boldsymbol{{0}}}}$ $\leq K_{0}$ by (A2),
@@
$\Upsilon_{\boldsymbol{\nu}}$ $\leq C_{R}\sum_{\boldsymbol{0}\neq\boldsymbol{m}\leq\boldsymbol{\nu}}\begin{pmatrix}\boldsymbol{\nu}\\ \boldsymbol{m}\end{pmatrix}\left\|\frac{\partial^{\boldsymbol{m}}_{\boldsymbol{y}}a(\boldsymbol{y})}{a(\boldsymbol{y})}\right\|_{L^{\infty}(D)}\Upsilon_{\boldsymbol{\nu}-\boldsymbol{m}}$ by (A3),
$\leq C_{R}C_{G}\sum_{\boldsymbol{0}\neq\boldsymbol{m}\leq\boldsymbol{\nu}}\begin{pmatrix}\boldsymbol{\nu}\\ \boldsymbol{m}\end{pmatrix}(|{\boldsymbol{{m}}}|!)^{\sigma}{{\boldsymbol{{b}}}}^{{\boldsymbol{{\nu}}}}\Upsilon_{\boldsymbol{\nu}-\boldsymbol{m}}$ by (A5),
@@ 3.2 Gevrey regular model on a bounded parameter domain
$a_{\max}=\sup_{\boldsymbol{x}\in D,\boldsymbol{y}\in\Omega}[a(\boldsymbol{y})](\boldsymbol{x})\qquad\text{ and }\qquad a_{\min}=\inf_{\boldsymbol{x}\in D,\boldsymbol{y}\in\Omega}[a(\boldsymbol{y})](\boldsymbol{x}).$
@@
Let assumptions (A1)(A5) and the assumptions of Theorem 2.1 hold with $\mu({\rm d}\boldsymbol{y})=\bigotimes_{j\geq 1}{\rm d}y_{j}$ and $U=[-\frac{1}{2},\frac{1}{2}]^{\mathbb{N}}$. Then there exists a generating vector constructed by the CBC algorithm such that for $\sigma p<1$,
$\sqrt{{\mathbb{E}}\big|I_{s}^{\varphi}(J(\boldsymbol{q}_{h,s}))-{Q_{{\boldsymbol{{\Delta}}}}^{\varphi}}(J(\boldsymbol{q}_{h,s}))\big|^{2}}=\begin{cases}\mathcal{O}(n^{-r/p+1/2})&\text{if}~p\in(\frac{2r}{3},\frac{1}{\sigma}),\\ \mathcal{O}(n^{-1+\varepsilon})\;\forall\varepsilon\in(0,\frac{2\sigma r-3}{2\sigma r-4})&\text{if}~p\in(0,\frac{2r}{3}],\end{cases}$
where the implied coefficient is independent of $s$ when the generating vector is obtained by a CBC algorithm using the weights
@@
Although the weights (19) are of a general form, an alternative construction given in [25] demonstrates how to obtain analogous weights in smoothness-driven product-and-order dependent (SPOD) form guaranteeing dimension-independent QMC convergence rates. The same holds for those weights in Theorem 3.12.
Proof 3.10 (Proof of Theorem 3.8).
@@
$\leq\frac{2^{|\mathfrak{u}|}C^{2r}_{E}C_{J}^{2r}}{a^{r}_{\min}}\sum_{\mathfrak{v}\subseteq\mathfrak{u}}\int_{[-1/2,1/2]^{s-|\mathfrak{u}|}}\bigg\|\frac{\partial^{|\mathfrak{v}|}}{\partial\boldsymbol{y}_{\mathfrak{v}}}\boldsymbol{q}_{h}(\cdot,\boldsymbol{y})\bigg\|_{Q_{h}}^{2r}\;\textup{d}\boldsymbol{y}_{-\mathfrak{u}}.$ by (A1),
@@
since (A2), (A3) and (A5) hold. Here, we use the notation $\boldsymbol{b}_{\mathfrak{u}}=\prod_{j\in\mathfrak{u}}b_{j}$ (note that for each subset $\mathfrak{u}\subset\{1,\ldots,s\}$, we can associate a multi-index $\boldsymbol{\nu}\in\mathbb{N}_{0}^{s}$ with $\nu_{j}=1$ for every $j\in\mathfrak{u}$ and $\nu_{j}=0$ otherwise). Then, setting $C^{2r}=\frac{C^{2r}_{E}C_{J}^{2r}}{a^{r}_{\min}}C^{2r}_{S}a^{r}_{\max}\|f\|^{2r}_{L^{2}(D)}$, and using the value of $C_{s,\gamma,\lambda}$ in (9),
@@
where $\Xi_{j}=\max\big\{1,2^{1/(2r)}\big(\frac{2\zeta(2\lambda)}{(2\pi^{2})^{\lambda}}\big)^{1/(2r\lambda)}\big\}(C_{R}C_{G}+1)b_{j}$. The first inequality is Jensen’s inequality in the version of [23, Thm. 19], and the second inequality follows from the definition of the $\Xi_{j}$. Noting that
@@
Here, we use an argument from [11]: We recast this sum as a sum over subsets as
@@
$\sum_{\boldsymbol{\nu}\in\{0:1\}^{s}}\bigg((|\boldsymbol{\nu}|!)^{\sigma r}\prod_{j=1}^{s}\Psi_{j}^{r\nu_{j}}\bigg)^{\frac{2\lambda}{1+\lambda}}\leq\sum_{\ell=0}^{\infty}(\ell!)^{\frac{2\lambda\sigma r}{1+\lambda}-1}\bigg(\sum_{j\geq 1}\Psi_{j}^{\frac{2\lambda r}{1+\lambda}}\bigg)^{\ell}$
@@
In the uniform and affine model (cf., e.g., [7, 11, 13, 15, 40, 31, 32, 39]), the uncertain coefficient is endowed with the parameterization
@@ 3.3 Unbounded and Gevrey regular model
Lemmas 3.5 and 3.4 with $K_{0}=C_{S}\sqrt{\overline{a}(\boldsymbol{y})}\|f\|_{L^{2}(D)}$, $K_{1}=C_{M}$, and $\Upsilon_{\boldsymbol{\nu}}=\|\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{Q_{h}}$ yield
$\|\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{Q_{h}}\leq C_{S}\sqrt{\overline{a}(\boldsymbol{y})}\|f\|_{{\mathcal{E}_{h}}}(C_{M}+1)^{|\boldsymbol{\nu}|}(|\boldsymbol{\nu}|!)^{\sigma}\boldsymbol{b}^{\boldsymbol{\nu}}.$
This bound can be used to control the weighted Sobolev norm (11) of $\|\boldsymbol{q}_{h}(\boldsymbol{y})\|_{Q_{h}}$.
@@
    $\|a(\boldsymbol{y})\|_{L^{\infty}(D)}^{r}\;\|a^{-1}(\boldsymbol{y})\|_{L^{\infty}(D)}^{r}\leq C_{Q}\prod_{j=1}^{\infty}\exp\left(2b_{j}|y_{j}|\right)$
    with the $b_{j}$ as characterized in (A5) and $r$ as in (A4).
@@
Let assumptions (A1)(A4) and the assumptions of Theorem 2.2 hold with $\mu=\bigotimes_{j\geq 1}\mathcal{N}(0,1)$ and $U=\mathbb{R}^{\mathbb{N}}$. Then, there exists a generating vector constructed by the CBC algorithm such that for $\sigma p<1$,
$\sqrt{{\mathbb{E}}\big|I_{s}^{\varphi}(J(\boldsymbol{q}_{h,s}))-{Q_{{\boldsymbol{{\Delta}}}}^{\varphi}}(J(\boldsymbol{q}_{h,s}))\big|^{2}}=\begin{cases}\mathcal{O}(n^{-r/p+1/2})&\text{if}~p\in(\frac{2r}{3},\frac{1}{\sigma}),\\ \mathcal{O}(n^{-1+\varepsilon})\;\forall\varepsilon\in(0,\frac{2\sigma r-3}{2\sigma r-4})&\text{if}~p\in(0,\frac{2r}{3}],\end{cases}$
where the implied coefficient is independent of $s$ when the generating vector is obtained by a CBC algorithm using the weights
@@
$\leq C_{J}^{2r}\int_{\mathbb{R}^{s-|\mathfrak{u}|}}\bigg(\sum_{\mathfrak{v}\subseteq\mathfrak{u}}\bigg\|\frac{\partial^{|\mathfrak{v}|}}{\partial\boldsymbol{y}_{\mathfrak{v}}}\boldsymbol{q}_{h}(\boldsymbol{y})\bigg\|_{L^{2}(D)}^{r}\bigg)^{2}\prod_{j\in\{1:s\}\setminus\mathfrak{u}}\varphi(y_{j})\,{\rm d}\boldsymbol{y}_{-\mathfrak{u}}$ by (A4)
$\leq C_{J}^{2r}C_{E}^{2r}\int_{\mathbb{R}^{s-|\mathfrak{u}|}}\frac{1}{{\underline{a}(\boldsymbol{y})^{r}}}\bigg(\sum_{\mathfrak{v}\subseteq\mathfrak{u}}\bigg\|\frac{\partial^{|\mathfrak{v}|}}{\partial\boldsymbol{y}_{\mathfrak{v}}}\boldsymbol{q}_{h}(\boldsymbol{y})\bigg\|_{Q_{h}}^{r}\bigg)^{2}\prod_{j\in\{1:s\}\setminus\mathfrak{u}}\varphi(y_{j})\,{\rm d}\boldsymbol{y}_{-\mathfrak{u}}$ by (A1),
$\leq 2^{|\mathfrak{u}|}C_{J}^{2r}C_{E}^{2r}\int_{\mathbb{R}^{s-|\mathfrak{u}|}}\frac{1}{{\underline{a}(\boldsymbol{y})^{r}}}\sum_{\mathfrak{v}\subseteq\mathfrak{u}}\bigg\|\frac{\partial^{|\mathfrak{v}|}}{\partial\boldsymbol{y}_{\mathfrak{v}}}\boldsymbol{q}_{h}(\boldsymbol{y})\bigg\|_{Q_{h}}^{2r}\prod_{j\in\{1:s\}\setminus\mathfrak{u}}\varphi(y_{j})\,{\rm d}\boldsymbol{y}_{-\mathfrak{u}}$ by Young’s ineq.
@@
Using (A4), Fubini’s theorem, and the integral identity $\int_{\mathbb{R}}{\rm e}^{2b_{j}|y|}\varphi(y)\,{\rm d}y=2{\rm e}^{2b_{j}^{2}}\Phi(2b_{j})$, we receive
@@
$\leq C_{Q}\int_{\mathbb{R}^{s}}\bigg(\prod_{j\in\{1:s\}\setminus\mathfrak{u}}{\rm e}^{2b_{j}|y_{j}|}\varphi(y_{j})\bigg)\bigg(\prod_{j\in\mathfrak{u}}{\rm e}^{2b_{j}|y_{j}|}\varpi_{j}^{2}(y_{j})\bigg)\,{\rm d}\boldsymbol{y}$ by (A4)
@@
In the lognormal case (cf., e.g., [16, 17, 18, 19, 24, 30, 41]), we define

### v2

@@ 3.1 The setting
If $J$ is a continuous linear functional then 4 holds with $r=1$, since we have $|\tfrac{\partial}{\partial\boldsymbol{q}_{h}}J(\boldsymbol{q}_{h})|\leq C_{J}$ for all $\boldsymbol{q}_{h}$ some constant $C_{J}$, and the chain rule dictates that
@@
which is an even stronger estimate than 4 with $r=1$.
@@
then 4 holds. This can be seen using the Leibniz rule for the product of two functions $f({\boldsymbol{{y}}})$ and $g({\boldsymbol{{y}}})$, whereby
@@
where we used the Cauchy–Schwarz inequality twice. This verifies 4 with $r=2$.
The next important assumption is on the random parameter. This assumption is motivated by [ChernovL24, ChernovL24b]. By Pringsheim’s theorem (cf., e.g., [pringsheim, p. 169]), the case $\sigma=1$ in (16) below corresponds to the class of holomorphic functions; the cases of $\sigma\in[0,1)$ may therefore be viewed in some abstract way as slightly smoother than holomorphic functions. Since $\sigma=1$ already recovers the theoretically expected rate, estimating $1\leq(|\boldsymbol{\nu}|!)^{\sigma}\leq|\boldsymbol{\nu}|!$ for $0\leq\sigma<1$ is effectively lossless. Thus, we avoid including the case $\sigma\in[0,1)$ in the analysis since some later inequalities exploit the fact that $x\mapsto x^{\sigma}$, $\sigma\geq 1$, is convex for $x\geq 0$. The assumed cases of $\sigma\geq 1$ covers holomorphic and less regular parametric dependence.
The Gevrey class consists of functions with a growth condition placed on their higher-order partial derivatives. However, in the unbounded setting, it is typically desirable to let the constant factor of the parametric regularity bound on $\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}a(\boldsymbol{y})\|_{L^{\infty}(D)}$ depend on the parametric variable $\boldsymbol{y}$ in order to be able to cover, e.g., lognormal random fields. To this end, we follow [GuKa24] and place the Gevrey assumption on the ratio $\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}a(\boldsymbol{y})/a(\boldsymbol{y})\|_{L^{\infty}(D)}$.
Assumption 5.
@@
Remark 3.3.
Note that Assumption (16) is stated in the non-dimensionally truncated parametric domain $U$.
Example 3.4.
Assumption 5 is satisfied by the classic uniform and affine expansion
$a(\boldsymbol{x},\boldsymbol{y})=a_{0}(\boldsymbol{x})+\sum_{j\in\mathbb{N}}y_{j}\psi_{j}(\boldsymbol{x}),\quad\boldsymbol{y}\in[-\tfrac{1}{2},\tfrac{1}{2}]^{\mathbb{N}},$
where $a_{0}\in L^{\infty}(D)$, $\psi_{j}\in L^{\infty}(D)$, and $(\|\psi_{j}\|_{L^{\infty}(D)})_{j\in\mathbb{N}}\in\ell^{1}(\mathbb{N})$. By exploiting the fact that
$\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}a(\boldsymbol{x},\boldsymbol{y})=\begin{cases}a(\boldsymbol{x},\boldsymbol{y})&\text{if}~\boldsymbol{\nu}=\mathbf{0}\\ \psi_{j}(\boldsymbol{x})&\text{if}~\boldsymbol{\nu}=\boldsymbol{e}_{j}\\ 0&\text{otherwise}\end{cases}$
there especially holds
$|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}a(\boldsymbol{x},\boldsymbol{y})|\leq\prod_{j\in{\rm supp}(\boldsymbol{\nu})}\|\psi_{j}\|_{L^{\infty}(D)}.$
Dividing through by $a(\boldsymbol{x},\boldsymbol{y})$ yields
$\bigg|\frac{\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}a(\boldsymbol{x},\boldsymbol{y})}{a(\boldsymbol{x},\boldsymbol{y})}\bigg|\leq\prod_{j\in{\rm supp}(\boldsymbol{\nu})}b_{j},\quad\text{where}~b_{j}=\frac{\|\psi_{j}\|_{L^{\infty}(D)}}{a_{\min}}.$
Since the higher-order partial derivatives of $a$ vanish for $\boldsymbol{\nu}\not\leq\mathbf{1}$, we can easily extend the bound for higher-order derivatives as
$\bigg\|\frac{\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}a(\cdot,\boldsymbol{y})}{a(\cdot,\boldsymbol{y})}\bigg\|_{L^{\infty}(D)}\leq(|\boldsymbol{\nu}|!)^{\sigma}\boldsymbol{b}^{\boldsymbol{\nu}},$
where $\sigma\geq 1$ is arbitrary.
@@
$a(\boldsymbol{x},\boldsymbol{y})=\exp\left(\sum_{j\in{\mathbb{N}}}{\xi(y_{j})\psi_{j}(\boldsymbol{x})}\right),$
where $\psi_{j}\in L^{\infty}(D)$ are as in Example 3.4. If we choose $\xi$ as the identity, we obtain the (classical, analytic) lognormal case, and for the (non-analytic) lognormal Gevrey-$\sigma$ case, we choose
@@
To complete this subsection, we point out a consequence of Assumptions 2, 3 and 5 (namely Lemma 3.7 below) that we shall use in our analysis. We also recall two other results we need whose proofs can be found elsewhere.
Lemma 3.6 (see [GuKa24, Lemma 3.1]).
@@
2, 3 and 5 imply
(17) $\|\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{\mathcal{W}_{h}}\leq C_{S}\|a({\boldsymbol{{y}}})\|^{1/2}_{L^{\infty}(D)}\|f\|_{L^{2}(D)}(C_{R}C_{G}+1)^{|\boldsymbol{\nu}|}(|\boldsymbol{\nu}|!)^{\sigma}\boldsymbol{b}^{\boldsymbol{\nu}}.$
Proof 3.8.
Set $\Upsilon_{\boldsymbol{\nu}}=\|\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{\mathcal{W}_{h}}$. Then
$\Upsilon_{{\boldsymbol{{0}}}}$ $\leq K_{0}$ by 2,
@@
$\Upsilon_{\boldsymbol{\nu}}$ $\leq C_{R}\sum_{\boldsymbol{0}\neq\boldsymbol{m}\leq\boldsymbol{\nu}}\begin{pmatrix}\boldsymbol{\nu}\\ \boldsymbol{m}\end{pmatrix}\left\|\frac{\partial^{\boldsymbol{m}}_{\boldsymbol{y}}a(\boldsymbol{y})}{a(\boldsymbol{y})}\right\|_{L^{\infty}(D)}\Upsilon_{\boldsymbol{\nu}-\boldsymbol{m}}$ by 3,
$\leq C_{R}C_{G}\sum_{\boldsymbol{0}\neq\boldsymbol{m}\leq\boldsymbol{\nu}}\begin{pmatrix}\boldsymbol{\nu}\\ \boldsymbol{m}\end{pmatrix}(|{\boldsymbol{{m}}}|!)^{\sigma}{{\boldsymbol{{b}}}}^{{\boldsymbol{{\nu}}}}\Upsilon_{\boldsymbol{\nu}-\boldsymbol{m}}$ by 5,
@@ 3.2 Gevrey regular model on a bounded parameter domain
$a_{\max}=\sup_{\boldsymbol{x}\in D,\,\boldsymbol{y}\in{U}}[a(\boldsymbol{y})](\boldsymbol{x})\qquad\text{ and }\qquad a_{\min}=\inf_{\boldsymbol{x}\in D,\,\boldsymbol{y}\in{U}}[a(\boldsymbol{y})](\boldsymbol{x}).$
@@
Let Assumptions 15 and the assumptions of Theorem 2.1 hold with $\mu({\rm d}\boldsymbol{y})=\bigotimes_{j\geq 1}{\rm d}y_{j}$ and $U=[-\frac{1}{2},\frac{1}{2}]^{\mathbb{N}}$. Then there exists a generating vector constructed by the CBC algorithm such that for $\sigma p<1$,
$\sqrt{{\mathbb{E}}\big|I_{s}^{\varphi}(J({\boldsymbol{q}_{h}}))-{Q_{{\boldsymbol{{\Delta}}}}^{\varphi}}(J({\boldsymbol{q}_{h}}))\big|^{2}}=\begin{cases}\mathcal{O}(n^{-r/p+1/2})&\text{if}~p\in(\frac{2r}{3},\frac{1}{\sigma}),\\ \mathcal{O}(n^{-1+\varepsilon})\;\forall\varepsilon\in(0,\frac{2\sigma r-3}{2\sigma r-4})&\text{if}~p\in(0,\frac{2r}{3}],\end{cases}$
where the implied coefficient, which depends on $\varepsilon$, is independent of $s$ when the generating vector is obtained by a CBC algorithm using the weights
@@
Although the weights (19) are of a general form, an alternative construction given in [KaarniojaKKNS22] demonstrates how to obtain analogous weights in smoothness-driven product-and-order dependent (SPOD) form guaranteeing dimension-independent QMC convergence rates. The same holds for those weights in Theorem 3.14.
Proof 3.12 (Proof of Theorem 3.10).
@@
$\leq\frac{2^{|\mathfrak{u}|}C^{2r}_{E}C_{J}^{2r}}{a^{r}_{\min}}\sum_{\mathfrak{v}\subseteq\mathfrak{u}}\int_{[-1/2,1/2]^{s-|\mathfrak{u}|}}\bigg\|\frac{\partial^{|\mathfrak{v}|}}{\partial\boldsymbol{y}_{\mathfrak{v}}}\boldsymbol{q}_{h}(\cdot,\boldsymbol{y})\bigg\|_{\mathcal{W}_{h}}^{2r}\;\textup{d}\boldsymbol{y}_{-\mathfrak{u}}.$ by 1,
@@
since 2, 3 and 5 hold. Here, we use the notation $\boldsymbol{b}_{\mathfrak{u}}=\prod_{j\in\mathfrak{u}}b_{j}$ (note that for each subset $\mathfrak{u}\subset\{1,\ldots,s\}$, we can associate a multi-index $\boldsymbol{\nu}\in\mathbb{N}_{0}^{s}$ with $\nu_{j}=1$ for every $j\in\mathfrak{u}$ and $\nu_{j}=0$ otherwise). Then, setting $C^{2r}=\frac{C^{2r}_{E}C_{J}^{2r}}{a^{r}_{\min}}C^{2r}_{S}a^{r}_{\max}\|f\|^{2r}_{L^{2}(D)}$, and using the value of $C_{s,\gamma,\lambda}$ in (9),
@@
where $\Xi_{j}=\max\big\{1,2^{1/(2r)}\big(\frac{2\zeta(2\lambda)}{(2\pi^{2})^{\lambda}}\big)^{1/(2r\lambda)}\big\}(C_{R}C_{G}+1)b_{j}$. The first inequality is Jensen’s inequality in the version of [HardyLP53, Thm. 19], and the second inequality follows from the definition of the $\Xi_{j}$. Noting that
@@
Here, we use an argument from [spodpaper14]: We recast this sum as a sum over subsets as
@@
$\sum_{\boldsymbol{\nu}\in\{0:1\}^{s}}\bigg((|\boldsymbol{\nu}|!)^{\sigma r}\prod_{j=1}^{s}\Psi_{j}^{r\nu_{j}}\bigg)^{\frac{2\lambda}{1+\lambda}}\leq\sum_{\ell=0}^{\infty}(\ell!)^{\frac{2\lambda\sigma r}{1+\lambda}-1}\bigg(\sum_{j\geq 1}\Psi_{j}^{\frac{2\lambda r}{1+\lambda}}\bigg)^{\ell}{.}$
@@
In the uniform and affine model (cf., e.g., [cds10, spodpaper14, dicklegiaschwab, ghs18, SchwabG11, kss12, kssmultilevel, schwab13]), the uncertain coefficient is endowed with the parameterization
@@ 3.3 Unbounded and Gevrey regular model
Lemmas 3.7 and 3.6 with $K_{0}=C_{S}\sqrt{\overline{a}(\boldsymbol{y})}\|f\|_{L^{2}(D)}$, $K_{1}=C_{M}$, and $\Upsilon_{\boldsymbol{\nu}}=\|\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{\mathcal{W}_{h}}$ yield
$\|\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{\mathcal{W}_{h}}\leq C_{S}\sqrt{\overline{a}(\boldsymbol{y})}\|f\|_{{\mathcal{E}_{h}}}(C_{M}+1)^{|\boldsymbol{\nu}|}(|\boldsymbol{\nu}|!)^{\sigma}\boldsymbol{b}^{\boldsymbol{\nu}}.$
This bound can be used to control the weighted Sobolev norm (11) of $\|\boldsymbol{q}_{h}(\boldsymbol{y})\|_{\mathcal{W}_{h}}$.
@@
$\|a(\boldsymbol{y})\|_{L^{\infty}(D)}^{r}\;\|a^{-1}(\boldsymbol{y})\|_{L^{\infty}(D)}^{r}\leq C_{Q}\prod_{j=1}^{\infty}\exp\left(2b_{j}|y_{j}|\right)\quad{\text{for all}~\boldsymbol{y}\in U}$
with the $b_{j}$ as characterized in 5 and $r$ as in 4.
Notably, while 5 is used to establish dimension-independent QMC convergence rates, 6 is needed to ensure that the expected value of the PDE response does not diverge. We remark that Assumption 6 needs to be enforced in the infinite-dimensional parameter space $U$.
@@
Let Assumptions 16 hold with $\mu=\bigotimes_{j\geq 1}\mathcal{N}(0,1)$ and $U=\mathbb{R}^{\mathbb{N}}$. Then, there exists a generating vector constructed by the CBC algorithm such that for $\sigma p<1$,
$\sqrt{{\mathbb{E}}\big|I_{s}^{\varphi}(J({\boldsymbol{q}_{h}}))-{Q_{{\boldsymbol{{\Delta}}}}^{\varphi}}(J({\boldsymbol{q}_{h}}))\big|^{2}}=\begin{cases}\mathcal{O}(n^{-r/p+1/2})&\text{if}~p\in(\frac{2r}{3},\frac{1}{\sigma}),\\ \mathcal{O}(n^{-1+\varepsilon})\;\forall\varepsilon\in(0,\frac{2\sigma r-3}{2\sigma r-4})&\text{if}~p\in(0,\frac{2r}{3}],\end{cases}$
where the implied coefficient, which depends on $\varepsilon$, is independent of $s$ when the generating vector is obtained by a CBC algorithm using the weights
@@
$\leq C_{J}^{2r}\int_{\mathbb{R}^{s-|\mathfrak{u}|}}\bigg(\sum_{\mathfrak{v}\subseteq\mathfrak{u}}\bigg\|\frac{\partial^{|\mathfrak{v}|}}{\partial\boldsymbol{y}_{\mathfrak{v}}}\boldsymbol{q}_{h}(\boldsymbol{y})\bigg\|_{L^{2}(D)}^{r}\bigg)^{2}\prod_{j\in\{1:s\}\setminus\mathfrak{u}}\varphi(y_{j})\,{\rm d}\boldsymbol{y}_{-\mathfrak{u}}$ by 4
$\leq C_{J}^{2r}C_{E}^{2r}\int_{\mathbb{R}^{s-|\mathfrak{u}|}}\frac{1}{{\underline{a}(\boldsymbol{y})^{r}}}\bigg(\sum_{\mathfrak{v}\subseteq\mathfrak{u}}\bigg\|\frac{\partial^{|\mathfrak{v}|}}{\partial\boldsymbol{y}_{\mathfrak{v}}}\boldsymbol{q}_{h}(\boldsymbol{y})\bigg\|_{\mathcal{W}_{h}}^{r}\bigg)^{2}\prod_{j\in\{1:s\}\setminus\mathfrak{u}}\varphi(y_{j})\,{\rm d}\boldsymbol{y}_{-\mathfrak{u}}$ by 1,
$\leq 2^{|\mathfrak{u}|}C_{J}^{2r}C_{E}^{2r}\int_{\mathbb{R}^{s-|\mathfrak{u}|}}\frac{1}{{\underline{a}(\boldsymbol{y})^{r}}}\sum_{\mathfrak{v}\subseteq\mathfrak{u}}\bigg\|\frac{\partial^{|\mathfrak{v}|}}{\partial\boldsymbol{y}_{\mathfrak{v}}}\boldsymbol{q}_{h}(\boldsymbol{y})\bigg\|_{\mathcal{W}_{h}}^{2r}\prod_{j\in\{1:s\}\setminus\mathfrak{u}}\varphi(y_{j})\,{\rm d}\boldsymbol{y}_{-\mathfrak{u}}$ by Young’s ineq.
@@
Using 6, Fubini’s theorem, and the integral identity $\int_{\mathbb{R}}{\rm e}^{2b_{j}|y|}\varphi(y)\,{\rm d}y=2{\rm e}^{2b_{j}^{2}}\Phi(2b_{j})$, we receive
@@
$\leq C_{Q}\int_{\mathbb{R}^{s}}\bigg(\prod_{j\in\{1:s\}\setminus\mathfrak{u}}{\rm e}^{2b_{j}|y_{j}|}\varphi(y_{j})\bigg)\bigg(\prod_{j\in\mathfrak{u}}{\rm e}^{2b_{j}|y_{j}|}\varpi_{j}^{2}(y_{j})\bigg)\,{\rm d}\boldsymbol{y}$ by 6
@@
In the lognormal case (cf., e.g., [gittelson, GrahamKNSSS15, GrahamKNSS11, log3, log4, log5, schwabtodor]), we define